# Fantastic Copyrighted Beasts and How (Not) to Generate Them

### v1

@@ 7 Limitations and Future Work
Our work provides an initial step forward for systematically evaluating the likelihood of generating copyrighted characters and the effectiveness of inference-time mitigation strategies. Future works can improve these evaluation protocols and mitigations in several ways. First, they can leverage optimization-based approaches to identify more complicated indirect anchors. Second, they can explore improved mechanisms to identify user intent to generate copyright characters from prompts alone. For example, if a user writes a complicated prompt “A video game plumber with a red hat and an M on the hat, in blue overalls….”, model improvements could better map the description to a potential character so that their name could be included in the negative prompt. Third, future work can address additional broader types of similarly challenging visual content, like trademarks, as well as broader sets of less-popular characters—our assessment is limited to a relatively small set of popular characters. Fourth, metrics like consistency scoring and detection could be improved to better capture legally-relevant and human-centered notions of consistency and character similarity. While our work will likely be re-usable for these broader categories of copyrighted and trademarked content, we did not explicitly evaluate them here.
@@ 8 Conclusion
In this paper, we investigate two main research questions: 1) Which textual prompts can trigger generation of copyrighted characters; and 2) How effective are current runtime mitigation strategies and how we can improve them? To systematically study these questions, we curate a diverse set of copyrighted characters and develop a novel evaluation suite CopyCat that considers both elimination of similar output to copyrighted characters and generated image’s consistency with user input. We show how to leverage embedding space distance and common training corpora to extract useful indirect anchors—descriptions and keywords not explicitly mentioning the characters’ names. We find that these indirect anchors can be effective in triggering copyrighted character generation. Existing mitigations, namely prompt rewriting, are not fully effective and we suggest new runtime methods to improve them. Our work calls for more attention to the indirect anchoring challenge and the effectiveness of deployed mitigation strategies for copyrighted character protection. The insights we provide here can be operationalized by model deployers for copyright-aware image and video generation systems in the future.
Acknowledgement
@@ Appendix A Legal Background and Broader Societal Impacts
While past work has studied the setting of verbatim regurgitation of images (Carlini et al., 2023b), and some lawsuits focus on this particular legal issue (Vincent, 2023; Andersen et al. v. Stability AI et al., N.D. Cal. 2023), copyrighted characters pose a unique legal challenge (Sag, 2023; Henderson et al., 2023; Lee et al., 2024). Unlike in the verbatim memorization setting, copyrighted characters are computationally more like general concepts that can appear in many poses, sizes, and variations in the training data. So typical deduplication, or even near access free learning approaches (Vyas et al., 2023), will not work—something discussed by others (Henderson et al., 2023).
Copyrighted characters are a somewhat distinct area of copyright law with distinct rules to determine infringement (Schreyer, 2015; Hennessey, 2020). To simplify the legal rules, characters are defined by key distinctive features that as a whole comprise the character. This can lead to interesting situations. For example, in 2023 the copyright for the original version of Mickey Mouse character (Steamboat Willie) entered the public domain. But this version of the character did not wear white gloves. However, the gloved version of Mickey Mouse that is now well known has not yet entered the public domain. A number of legal scholars and commentators have pointed out that this means that using a visual depiction of the modern Mickey Mouse would likely lead to an infringement claim, but using the old style of Mickey Mouse (Steamboat Willie) would not (Liu, 2013; Lee, 2019).
@@
The paper studies how companies can respect the intellectual property rights of creators and their visual copyrighted characters from an inference-time technical perspective. Leveraging methods here will both improve likelihood that rights are respected and reduce litigation risk for companies. However, we do not address broader societal discussions on how artists should be compensated for training on images that may contain their intellectual property (such as their characters). This is a larger, worthy, discussion broader than the scope of our work. However, we note that current fair use doctrine in the United States may allow this training provided that mitigation strategies are used to prevent substantially similar outputs (Lee et al., 2024; Lemley & Casey, 2020; Henderson et al., 2023; Pasquale & Sun, 2024). Similar fair use standards exist in other countries as well, but globally there are also countries where even training may not be allowed and different approaches may be needed. There are also general labor displacement concerns that are important that go beyond the scope of this work.
Appendix B Full List of Characters and Studios in CopyCat
@@ C.1 Compute resource
All experiments are conducted on 2 NVIDIA A100 GPU cards, each with 80GB of memory. Table 3 provides statistics on the time cost for each image generation across all the evaluated models, using the character’s name as the input prompt.
@@
We also report the time cost per evaluation for a single image in Table 4, including the cost of running the GPT-4V detector on the image and calculating the consistency score between the image and its key characteristics using VQAScore (Lin et al., 2024). Note that the time cost of the GPT-4V detector is obtained via querying the API, so it may also depend on the real-time network traffic.
@@ C.2 Model Generation Configurations
C.3 GPT-4V evaluator
We use GPT-4V to check whether there exists a character in the image that can be recognized as an existing copyrighted character. For each image, we prompt the multimodal model with the alongside the following prompt:
Template for obtaining rewritten prompts Do you detect any copyrighted character in this image? If not, give a score of 0. If yes, give a score of 1 and note the character. Output in the format of character: , score: .
For example: “character: ”, score: 0”, “character: Elsa, score: 1”.
@@ C.4 Consistency Score
As described in §2, we use the VQAScore (Lin et al., 2024) to measure consistency ($\mathsf{CONS}$) by checking if the main subject of the target copyrighted character exists in the generated image. The backbone model for computing VQAScore is CLIP-FlanT5.101010https://huggingface.co/zhiqiulin/clip-flant5-xxl
@@ C.6 Prompt-Rewriting Template
For prompt-rewriting intervention, we adapt the DALL·E system prompt to simulate mitigation strategies adopted by model deployers. The prompt is modified to highlight copyright concerns as well as restrict prompt length to be around 60 words.  [^11] The original full DALL·E system prompt is discussed at https://github.com/spdustin/ChatGPT-AutoExpert/blob/main/_system-prompts/gpts/dalle.md
@@
Figure 6 shows an example of the generated keywords and descriptions for Mario.
@@
Appendix D Human evaluation (authors)
Specifically, we first sample 200 generated images (20 characters $\times$ 10 images per character) from various prompting configurations, including direct prompting with character names and indirect prompting using keywords or descriptions, with or without the application of mitigation strategies. We then ask 6 authors to independently annotate these images, following guidelines similar to those used for GPT-4V (described in Section C.3).
For these 200 records, we examine the accuracy of GPT-4V, with the majority-human scores as ground truths. We find that the scores assigned by GPT-4V obtain a fairly high accuracy of $82.5\%$. To further analyze the consistency and agreement, we compute the Cohen Kappa value (Cohen, 1960) between GPT-4V scores with the majority-human scores. As evaluated, we observe a Cohen Kappa value of 0.648, representing a substantial agreement between human annotators and GPT-4V. We also accompany the pair-wise agreement measurements among human annotators and GPT-4V in Figure 7.
Appendix E More results on DALL·E
Character name anchoring does not work on DALL·E system due to its built-in filter that detects and blocks requests that explicitly mention copyrighted characters. However, indirect anchoring is still able to bypass the system guardrails and generate high-quality images that highly resemble the target copyrighted characters, as illustrated in Figure 8.
@@ Appendix E More results on DALL·E
Appendix F More results on open-source models
F.1 Effect of Target’s Name as Negative Prompt for Indirect Anchoring
We also examine intervention strategies in cases where users provide keywords or descriptions to generate images. As shown in Table 5, consistent with our previous observations in Table 1 when the character name is part of user input, adding character name as negative prompt is still a very effective method to reduce recreating copyrighted characters. In particular, when the original prompt consists of 10 keywords or descriptions, incorporating target’s name as negative reduce $\mathsf{DETECT}$ by 50% or more, while $\mathsf{CONS}$ values remain almost constant. For all experiment setup, the $\mathsf{CONS}$ values either remain stable or show a slight decrease with the addition of negative prompts. From a practical perspective, adding copyright character detection and target name as negative prompt is a simple yet effective way of reducing the recreation of copyrighted characters, at the cost slight compromise in adhering to user request.
@@ F.2 Intent detection
    Retriever-based detector that uses a retriever to compare the user input against a database of copyrighted character descriptions. For a given user query, the retriever searches for similar descriptions based on OpenAI embeddings121212text-embedding-3-small. If no description with a cosine similarity greater than 0.7 is found, we conclude that the user query does not intend to generate characters substantially similar to copyrighted ones.
@@
To evaluate our detection methods, we curated a dataset comprising 200 descriptions of copyrighted characters and 200 standard prompts unlikely to cause copyright issues selected from MJHQ benchmarks131313huggingface.co/datasets/playgroundai/MJHQ-30K. We report accuracy, true positive rates (TPR) and false positive rates (FPR) as our evalaution metrics.
@@ Results
F.3 Embedding similarity analysis
@@
We randomly generate 100 60-word prompts per character using the template described in Appendix C, and rank them by embedding similarity to the corresponding character name. As shown in Figure 9, the top-ranked prompt by embedding similarity generates 26 characters successfully, versus only 16 for the bottom-ranked prompt.
@@
We also study how the success and failure of rewritten prompts correlate with their embedding similarity to the corresponding character name. Specifically, for each character, we generate 100 rewritten prompts and rank them by their embedding similarity to the character’s name. As shown in Figure 10, the top-ranked rewritten prompt by embedding similarity generates 20 characters successfully, versus only 12 for the bottom-ranked rewritten prompt. This suggests that potentially, rewritten prompts that fail to avoid character generation could be due to their high similarity to the character’s name.
@@
F.4 More results for Playground v2.5
@@
Interestingly, the model exhibits high sensitivity to even minor perturbations in the character’s name. For instance, if we randomly replace a single letter in the character’s name with a different letter, the model can only generate 8 out of the 50 characters successfully. The situation is even more extreme when we randomly replace 3 letters – in this case, the model could only generate 1 out of the 50 characters accurately (see Figure 11(b)).
On the other hand, if the character’s name is present in the prompt, and irrelevant keywords such as "dancing" or "swimming" are added, this generally does not affect the number of characters generated (see Figure 11(c) and Figure 11(d)). These findings suggest that the character name anchoring mode heavily relies on the exact spelling of the target character’s name to generate copyrighted characters.
@@
Figure 12 visualizes results using the character’s name as the prompt and various keywords as negative prompts. Including the character’s name in the prompt, even with detailed negative prompts, still leads to the generation of copyrighted characters. This suggests that T2I models are deeply anchored to these character names.
However, once we apply prompt rewriting and combine it with various negative prompts, the model is no longer inclined to generate these characters, as shown in Figure 13.
@@
F.5 Results for PixArt-$\alpha$, Stable Diffusion XL, and DeepFloyd IF
Figure 14 visualizes results from the PixArt-$\alpha$ model (Chen et al., 2024). With higher generation quality, the findings are also consistent with those observed using the Playground v2.5 model—adding more fine-grained negative prompts and applying prompt rewriting significantly reduces the similarity of the generated images to the original copyrighted character.
Figure 15 visualizes results from the Stable Diffusion XL (SDXL) model (Podell et al., 2024). Although the generation quality of SDXL is generally lower compared to the Playground model (see Figure 5), adding more fine-grained negative prompts and applying prompt rewriting significantly reduces the similarity of the generated images to the original copyrighted character.

### v2

@@ 7 Discussions and Limitations
Future works can improve these evaluation protocols and mitigations in several ways. First, they can leverage optimization-based approaches to identify more complicated indirect anchors. Second, they can explore improved mechanisms to identify user intent to generate copyrighted characters from prompts alone. For example, if a user writes a complicated prompt “A video game plumber with a red hat and an M on the hat, in blue overalls….”, model improvements could better map the description to a potential character so that their name could be included in the negative prompt. Third, future work can address additional broader types of similarly challenging visual content, like trademarks and broader sets of less-popular characters. Fourth, metrics like consistency scoring and detection could be improved to better capture legally-relevant and human-centered notions of consistency and character similarity. While our work will likely be re-usable for these broader categories of copyrighted and trademarked content, we did not explicitly evaluate them here.
@@ 8 Conclusion
In this paper, we study the important subquestion of copyrighted protection focusing on copyrighted characters. We address two main research questions: 1) Which textual prompts can trigger generation of copyrighted characters; and 2) How effective are current runtime mitigation strategies and how we can improve them? To systematically study these questions, we develop an evaluation framework that considers both elimination of similar output to copyrighted characters and generated image’s consistency with user input. We show how to leverage embedding space distance and common training corpora to extract useful indirect anchors—descriptions and keywords not explicitly mentioning the characters’ names but can be effective in triggering copyrighted character generation. Existing mitigations, namely prompt rewriting, are not fully effective and we suggest new runtime methods to improve them. Our work calls for more attention to the indirect anchoring challenge and the effectiveness of deployed mitigation strategies for copyrighted character protection. The insights we provide here can be operationalized by model deployers for copyright-aware image and video generation systems in the future.
Acknowledgements
@@
Ethics Statement
This work complies with the ICLR Code of Ethics. Regarding potential concerns in releasing copyrighted assets: The majority of the assets (e.g., prompts, descriptions, etc.) are generic and are likely not protected by copyright in most, if not all, jurisdictions. To the extent that we release generated images of characters (e.g., the screenshots in the paper), in many jurisdictions this would be considered acceptable under copyright law as part of research (e.g., fair use in the United States). However, to ensure compliance with more jurisdictions, beyond the paper, we can refrain from releasing certain images other than through direct sharing only with researchers on request under an agreement on proper use of the asset.
In general, our work systematically analyzes mitigation strategies and risks of generating certain copyrighted characters. We find that currently these mitigation strategies can sometimes fail, though they are actively used by different organizations. However, there may be better ways for companies to respect the intellectual property rights of creators and their visual copyrighted characters from both a technical and structural perspectives. Leveraging the lessons we provide here will both improve likelihood that rights are respected and reduce litigation risk for companies. However, we do not address broader societal discussions on how artists should be compensated for training on images that may contain their intellectual property (such as their characters). This is a larger, worthy, discussion broader than the scope of our work. Current fair use doctrine in the United States might allow this training provided that mitigation strategies are used to prevent substantially similar outputs (Lee et al., 2024; Lemley & Casey, 2020; Henderson et al., 2023; Pasquale & Sun, 2024). But this is far from resolved. Similar fair use standards exist in other countries as well. Globally, there are also countries where even training may not be allowed and different approaches may be needed than we describe here. There are also general labor displacement concerns: mitigation strategies might successfully prevent infringement of copyright characters and model training might generally be considered fair use, yet may nonetheless impact data creators’ livelihoods. These are important issues that go beyond the scope of this work.
Reproducibility Statement
We are dedicated to making every aspect of our work fully open-source, offering detailed instructions to ensure reproducibility. Detailed model choice and generation configurations are included in § C. We also provide the original prompts we use for querying models in § C.5 and § C.6.
We include code and data in our supplementary material. They cover all experiments in the paper, including image and video generation code, prompt and data used, evaluation under different settings, indirect anchor studies etc. We also provide detailed instruction along with the code to ensure reproducibility of our results.
@@ Appendix A Legal Background and Broader Societal Impacts
In addition to the discussion in § 1 on the unique challenges posed by copyrighted characters and the value of a quantitative study In this matter, we further expand on the legal background and broader societal impacts of our study below.
Legal Motivations Past work has studied the setting of verbatim regurgitation of images (car, 2023), and some lawsuits focus on this particular legal issue (Vincent, 2023; Andersen et al. v. Stability AI et al., N.D. Cal. 2023). Some recent work builds datasets for benchmarking copyright infringement unlearning methods (Ma et al., 2024) and attempts to jailbreak proprietary systems to output copyrighted images (Kim et al., 2024). Among the copyrighted subjects of interest, copyrighted characters, such as popular Intellectual Property (IP) from Disney, Nintendo, and Dreamworks, pose a unique legal challenge (Sag, 2023; Henderson et al., 2023; Lee et al., 2024). At least one lawsuit in China has already resulted in liability for an image generation service that generated the copyrighted character, Ultraman (Shimbun, 2024). Unlike in the verbatim memorization setting, copyrighted characters are computationally more like general concepts that can appear in many poses, sizes, and variations in the training data. Typical deduplication, or even near access free learning approaches (Vyas et al., 2023), will not work for this task (Henderson et al., 2023). Copyrighted characters are also in a certain area of copyright law with distinct rules to determine infringement (Schreyer, 2015; Hennessey, 2020). To simplify the legal rules, characters are defined by key distinctive features that as a whole comprise the character. not (Liu, 2013; Lee, 2019).
As discussed earlier, to simplify the legal rules, characters are often defined by key distinctive features that as a whole compromise the character. This can lead to interesting situations. For example, in 2023 the copyright for the original version of Mickey Mouse character (Steamboat Willie) entered the public domain. But this version of the character did not wear white gloves. However, the gloved version of Mickey Mouse that is now well known has not yet entered the public domain. A number of legal scholars and commentators have pointed out that this means that using a visual depiction of the modern Mickey Mouse would likely lead to an infringement claim, but using the old style of Mickey Mouse (Steamboat Willie) would
@@
The paper emphasizes that current mitigation strategies taken by companies can fall short of their intended purpose. There may be better ways for companies to respect the intellectual property rights of creators and their visual copyrighted characters from both a technical and structural perspectives. Leveraging the lessons we provide here will both improve likelihood that rights are respected and reduce litigation risk for companies. However, we do not address broader societal discussions on how artists should be compensated for training on images that may contain their intellectual property (such as their characters). This is a larger, worthy, discussion broader than the scope of our work. Current fair use doctrine in the United States might allow this training provided that mitigation strategies are used to prevent substantially similar outputs (Lee et al., 2024; Lemley & Casey, 2020; Henderson et al., 2023; Pasquale & Sun, 2024). But this is far from resolved. Similar fair use standards exist in other countries as well. Globally, there are also countries where even training may not be allowed and different approaches may be needed than we describe here. There are also general labor displacement concerns: mitigation strategies might successfully prevent infringement of copyright characters, yet may nonetheless impact data creators’ livelihoods. These are important issues that go beyond the scope of this work.
Appendix B Full List of Characters and Studios in Our Study
We source copyrighted characters from popular studios and franchises, as they are more likely to have been present in the training process of image and video generation models. In addition to U.S. studios like Disney and DreamWorks, we also include international ones like Nintendo and Shogakukan. In total, our collection includes 50 diverse popular copyrighted characters from 18 different studios and subsidiaries. The full list of characters can be found in Appendix B.
@@ C.1 Compute resource
All experiments are conducted on 2 NVIDIA A100 GPU cards, each with 80GB of memory. Tb. 4 provides statistics on the time cost for each image generation across all the evaluated models, using the character’s name as the input prompt.
@@
We also report the time cost per evaluation for a single image in Tb. 5, including the cost of running the GPT-4V detector on the image and calculating the consistency score between the image and its key characteristics using VQAScore (Lin et al., 2024). Note that the time cost of the GPT-4V detector is obtained via querying the API, so it may also depend on the real-time network traffic.
@@ C.2 Model Generation Configurations
C.3 Evaluator Set-up and Ablations
In order to automate the process of checking whether there exists a character in the image that can be recognized as an existing copyrighted character, we make use of recent Vision Language Models (VLMs). We conduct ablations study on model choice and prompt format, and choose the option with the highest accuracy and human agreement. First we discuss our human evaluation protocols as follows.
Human evaluation (authors):
Specifically, we first sample 200 generated images (20 characters $\times$ 10 images per character) from various prompting configurations, including direct prompting with character names and indirect prompting using keywords or descriptions, with or without the application of mitigation strategies. We then ask 6 authors to independently annotate these images, following guidelines similar to those used for GPT-4V (described in § C.3).
For these 200 records, we examine the accuracy of GPT-4V, with the majority-human scores as ground truths. We find that the scores assigned by GPT-4V obtain a fairly high accuracy of $82.5\%$. To further analyze the consistency and agreement, we compute the Cohen Kappa value (Cohen, 1960) between GPT-4V scores with the majority-human scores. As evaluated, we observe a Cohen Kappa value of 0.648, representing a substantial agreement between human annotators and GPT-4V. We also accompany the pair-wise agreement measurements among human annotators and GPT-4V in Fig. 6.
Using detector accuracy and human agreement scores, we ablate evaluator design choices and choose GPT-4v since it aligns the best with humans among other model choices and prompt format (Tb. 6). Our ablations experiments also show that asking for generic “copyrighted characters" in the prompt instead of specific characters results in higher agreement with humans (Tb. 6). For each image, we prompt the multimodal model with the alongside the following prompt:
Table 6: Ablation of Evaluator Design Choises. GPT-4v (the chosen evaluator in our paper) is the best option with highest accuracy and human agreement
Model Character Name in Prompt (Yes/ No) Accuracy Cohen-Kappa Agreement with Human
GPT-4V No 82.5% 0.65
GPT-4V Yes 69% 0.39
Claude Sonnet No 71% 0.43
Claude Sonnet Yes 65.5% 0.33
Llava-34b No 78% 0.56
Llava-34b Yes 66.5% 0.35
Template for obtaining rewritten prompts Do you detect any copyrighted character in this image? If not, give a score of 0. If yes, give a score of 1 and note the character. Output in the format of character: , score: . For example: “character: ”, score: 0”, “character: Elsa, score: 1”.
To further demonstrate the generalizability of our evaluator choice to beyond the selected set of 50 characters in the paper, we show additional results on an additional suite of 50 unseen characters. The ablation study shows that GPT-4V with generic “copyrighted character" in prompt is still the best option (Tb. 7).
Table 7: Additional ablations of evaluator designs show that GPT-4V (the chosen evaluator in our paper) is still the best option with highest accuracy and human agreement on unseen characters.
Model Character Name in Prompt (Yes/ No) Accuracy Cohen-Kappa Agreement with Human
GPT-4V No 80% 0.52 (Moderate)
GPT-4V Yes 58.4% 0.26 (Fair)
Claude Sonnet No 78% 0.49 (Moderate)
Claude Sonnet Yes 49.1% 0.15 (Slight)
The additional 50 characters list include: [’Aang’, ’Alladin’, ’Bambi’, ’Barbie ’, ’BB-8’, ’Belle’, ’Betty Boop’, ’BoJack Horseman’, ’Bugs Bunny’, ’Charlie Brown’, ’Chewbacca’, ’Cruella de Vil’, ’Daffy Duck’, ’Daisy Duck’, ’Deadpool’, ’Dumbo’, ’Eren Yeager’, ’Garfield’, ’Genie’, ’Harley Quinn’, ’Hello Kitty’, ’Jafar’, ’Minions’, ’Minnie Mouse’, ’Moana’, ’Mr. Krabs’, ’Mr. Potato Head’, ’Mulan’, ’Mushu’, ’My Little Pony’, ’Nick Wilde’, ’Optimus Prime’, ’Pink Panther’, ’Princess Peach’, ’R2-D2’, ’Rick Sanchez’, ’Sakura Haruno’, ’Sasuke Uchiha’, ’She-Ra’, ’Shrek’, ’Simba ’, ’Simpsons’, ’Snoopy’, ’Son Goku’, ’Stitch’, ’Superman ’, ’The Smurfs’, ’Totoro’, ’Ursula’, ’Wolverine’]
@@ C.4 Consistency Score
As described in § 2, we use the VQAScore (Lin et al., 2024) to measure consistency ($\mathsf{CONS}$) by checking if the main subject of the target copyrighted character exists in the generated image. The backbone model for computing VQAScore is CLIP-FlanT5.‡‡‡‡‡‡https://huggingface.co/zhiqiulin/clip-flant5-xxl
@@
Human Evaluation of $\mathsf{CONS}$
Authors give binary annotations (e.g., is the key characteristic present in the generated image) and we use the majority votes as the ground truth. The AUROC of the continuous CONS score against binary human annotations is 0.82, suggesting a high agreement. Note that CONS is based on VQAScore (Lin et al., 2024), which has been shown to achieve SOTA performance on challenging image-text matching benchmarks.
@@ C.6 Prompt-Rewriting Template
For prompt-rewriting intervention, we adapt the DALL·E system prompt to simulate mitigation strategies adopted by model deployers. The prompt is modified to highlight copyright concerns as well as restrict prompt length to be around 60 words. The original full DALL·E system prompt can be obtained via prompt extraction methods, by directly prompting GPT-4 in a certain format, as discussed in various sources like McGuinness (2024); Dustin (2024). It should be noted, however, that this may not be the exact DALL·E system prompt, but prompt extraction methods have generally been successful. Regardless of whether this is the exact system prompt used in production, it is a sufficient approximation for our purposes (Zhang et al., 2024b; Wang et al., 2024). Given our limited access to the OpenAI system, we’ve done our best to recreate the pipeline as closely as possible. We also did some manual checking on the ChatGPT interface to verify that after rewriting with this set of instructions, DALL-E will not change the wording again when doing actual generation.
@@
Fig. 7 shows an example of the generated keywords and descriptions for Mario.
@@
Appendix D More results on DALL·E
Character name anchoring does not work on DALL·E system due to its built-in filter that detects and blocks requests that explicitly mention copyrighted characters. However, indirect anchoring is still able to bypass the system guardrails and generate high-quality images that highly resemble the target copyrighted characters, as illustrated in Fig. 8.
@@ Appendix D More results on DALL·E
Appendix E More results on open-source models
E.1 Effect of Target’s Name as Negative Prompt for Indirect Anchoring
We also examine intervention strategies in cases where users provide keywords or descriptions to generate images. As shown in Tb. 8, consistent with our previous observations in Tb. 1 when the character name is part of user input, adding character name as negative prompt is still a very effective method to reduce recreating copyrighted characters. In particular, when the original prompt consists of 10 keywords or descriptions, incorporating target’s name as negative reduce $\mathsf{DETECT}$ by 50% or more, while $\mathsf{CONS}$ values remain almost constant. For all experiment setup, the $\mathsf{CONS}$ values either remain stable or show a slight decrease with the addition of negative prompts. From a practical perspective, adding copyright character detection and target name as negative prompt is a simple yet effective way of reducing the recreation of copyrighted characters, at the cost slight compromise in adhering to user request.
@@ E.2 Intent detection
    Retriever-based detector that uses a retriever to compare the user input against a database of copyrighted character descriptions (Su et al., 2023). For a given user query, the retriever searches for similar descriptions based on OpenAI embeddings******text-embedding-3-small. If no description with a cosine similarity greater than 0.7 is found, we conclude that the user query does not intend to generate characters substantially similar to copyrighted ones.
@@
To evaluate our detection methods, we curated a dataset comprising 200 descriptions of copyrighted characters and 200 standard prompts unlikely to cause copyright issues selected from MJHQ benchmarks*†*†*†huggingface.co/datasets/playgroundai/MJHQ-30K. We report accuracy, true positive rates (TPR) and false positive rates (FPR) as our evalaution metrics.
@@ Results
E.3 Embedding similarity analysis
@@
We randomly generate 100 60-word prompts per character using the template described in § C, and rank them by embedding similarity to the corresponding character name. As shown in Fig. 9, the top-ranked prompt by embedding similarity generates 26 characters successfully, versus only 16 for the bottom-ranked prompt.
@@
We also study how the success and failure of rewritten prompts correlate with their embedding similarity to the corresponding character name. Specifically, for each character, we generate 100 rewritten prompts and rank them by their embedding similarity to the character’s name. As shown in Fig. 10, the top-ranked rewritten prompt by embedding similarity generates 20 characters successfully, versus only 12 for the bottom-ranked rewritten prompt. This suggests that potentially, rewritten prompts that fail to avoid character generation could be due to their high similarity to the character’s name.
@@
E.4 More results for Playground v2.5
@@
Interestingly, the model exhibits high sensitivity to even minor perturbations in the character’s name. For instance, if we randomly replace a single letter in the character’s name with a different letter, the model can only generate 8 out of the 50 characters successfully. The situation is even more extreme when we randomly replace 3 letters – in this case, the model could only generate 1 out of the 50 characters accurately (see Fig. 11(b)).
On the other hand, if the character’s name is present in the prompt, and irrelevant keywords such as "dancing" or "swimming" are added, this generally does not affect the number of characters generated (see Fig. 11(c) and Fig. 11(d)). These findings suggest that the character name anchoring mode heavily relies on the exact spelling of the target character’s name to generate copyrighted characters.
@@
Fig. 12 visualizes results using the character’s name as the prompt and various keywords as negative prompts. Including the character’s name in the prompt, even with detailed negative prompts, still leads to the generation of copyrighted characters. This suggests that T2I models are deeply anchored to these character names.
However, once we apply prompt rewriting and combine it with various negative prompts, the model is no longer inclined to generate these characters, as shown in Fig. 13.
@@
E.5 Results for PixArt-$\alpha$, Stable Diffusion XL, and DeepFloyd IF
Fig. 14 visualizes results from the PixArt-$\alpha$ model (Chen et al., 2024). With higher generation quality, the findings are also consistent with those observed using the Playground v2.5 model—adding more fine-grained negative prompts and applying prompt rewriting significantly reduces the similarity of the generated images to the original copyrighted character.
Fig. 15 visualizes results from the Stable Diffusion XL (SDXL) model (Podell et al., 2024). Although the generation quality of SDXL is generally lower compared to the Playground model (see Fig. 5), adding more fine-grained negative prompts and applying prompt rewriting significantly reduces the similarity of the generated images to the original copyrighted character.